OpenStudy (gabylovesyou):

Express 16 = 2^x as a logarithmic equation. log2x = 16 log162 = x log216 = x log16x = 2

OpenStudy (freckles):

\[y=b^x \implies \log_b(y)=x \\ \text{ where } b \in (0,1) \cup (1,\infty) , y>0\]

OpenStudy (gabylovesyou):

D ?

OpenStudy (gabylovesyou):

@freckles

OpenStudy (freckles):

can you compare y=b^x to your equation ?

OpenStudy (freckles):

and from it can you identify b and y?

OpenStudy (freckles):

and just plug in into the form that is equivalent that I have written above

OpenStudy (freckles):

base number is the one that ends up being the subscript of that log thing

OpenStudy (gabylovesyou):

mmm ;/

OpenStudy (agent0smith):

Look at what freckles gave you, y=b^x Then look at what you have, 16=2^x All you have to do is identify y and b, then plug them into log_b (y) = x

OpenStudy (gabylovesyou):

y = 16 b = 2 log_2 (16) = x

OpenStudy (gabylovesyou):

so C ?

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

It's hard to tell, but \[\Large 16 = 2^x\] should turn into \[\Large \log_2(16) = x\]

OpenStudy (agent0smith):

@Gabylovesyou I gave you the medal because you were correct.